# Sub-GeV millicharge dark matter from the $U(1)_{X}$ hidden sector

### v2

@@ Acknowledgments:
Appendix A Kinetic mixing and mass mixing in the Stueckelberg extension
@@
$\mathcal{R}=KO^{\prime}R_{1}^{\prime}R_{2}^{\prime}$ $=\begin{pmatrix}\frac{1}{\sqrt{1-\delta^{2}}}&0&0\\ -\frac{\delta}{\sqrt{1-\delta^{2}}}&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}\cos\phi&-\sin\phi&0\\ \sin\phi&\cos\phi&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\cos\psi&0&\sin\psi\\ 0&1&0\\ -\sin\psi&0&\cos\psi\end{pmatrix}$ (A.21)
@@ 1. $\chi\bar{\chi}\rightarrow\gamma,\gamma^{\prime},Z\rightarrow i\bar{i}$
$\sigma_{\chi\bar{\chi}\rightarrow i\bar{i}}(s)=\frac{g_{X}^{2}Q_{\chi}^{2}}{48\pi}\sqrt{\frac{s-4m_{i}^{2}}{s-4m_{\chi}^{2}}}\left(1+\frac{2m_{\chi}^{2}}{s}\right)\times$ (C.1)
@@
$+\frac{4(g_{2}s_{\theta}Q_{i}\mathcal{R}_{12})^{2}\left(2m_{i}^{2}+s\right)}{s^{2}}\bigg{\}}\,,$

### v3

@@ Acknowledgments:
Appendix A Kinetic and mass mixing in the Stueckelberg extension
@@
$\mathcal{R}=KO^{\prime}R_{1}^{\prime}R_{2}^{\prime}$ $=\begin{pmatrix}\frac{1}{\sqrt{1-\delta^{2}}}&0&0\\ -\frac{\delta}{\sqrt{1-\delta^{2}}}&1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}\cos\phi&-\sin\phi&0\\ \sin\phi&\cos\phi&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\cos\psi&0&\sin\psi\\ 0&1&0\\ -\sin\psi&0&\cos\psi\end{pmatrix}$
@@ 1. $\chi\bar{\chi}\rightarrow\gamma,\gamma^{\prime},Z\rightarrow i\bar{i}$
$\sigma_{\chi\bar{\chi}\rightarrow i\bar{i}}(s)$ $=\frac{g_{X}^{2}Q_{\chi}^{2}}{48\pi}\sqrt{\frac{s-4m_{i}^{2}}{s-4m_{\chi}^{2}}}\left(1+\frac{2m_{\chi}^{2}}{s}\right)\times$
@@
$+\frac{4(g_{2}s_{\theta}Q_{i}\mathcal{R}_{12})^{2}\left(2m_{i}^{2}+s\right)}{s^{2}}\bigg{\}}\,,$ (C.1)